OpenStudy (anonymous):

Which of the following systems of equations has no solution? a. 9x + 5y = 1; 15y = 18x − 4 b. -7x − 7 = 3y; -14y − 8 = -6x c. 4x − 3y = 9; 6y = 8x − 18 d. 7y = 5x − 10; 10x − 14y = 8

OpenStudy (anonymous):

Check for D..

OpenStudy (callisto):

Check the slope... if they have the same slope, then it has no solution since they don't intersect.

OpenStudy (anonymous):

5x - 7y = 10 Multiply it by 2, 10x - 14y = 20 10x - 14y = 8 On subtracting both x and y will get cancel and therefore no solution..

OpenStudy (callisto):

To find the slope, you can re-arrange the equations into the form: y=mx+c , slope = m or to rearrange the equations into the form of \(ax+by+c=0\), slope = -a/b